how do you solve this?
2(3h+2)-4h= -16

Answer by DrBeeee(684)

Solve
(1) 2(3h+2)-4h = -16
First step is to clear the (), yielding
(2) 6h + 4 -4h = -16
Next combine like terms and get
(3) 2h = -20
Now divide each side by 2 and get
(4) h = -20/2 or
(5) h = -10
Check this value using (1).
Is (2*(3*(-10)+2)-4(-10) = -16)?
Is (2*(-30+2)+40 = -16)?
Is (2*(-28)+40 = -16)?
Is (-56+40 = -16)?
Is ( -16 = -16)? Yes
Answer: h equals -10.